# Independent sets of non-geometric lattices and the maximal adjoint

### v2

@@ 4. Lattice operations
Let $\mathcal{L}$ be a finite, atomic, and graded lattice, and $A\in I(\mathcal{L})$ an independent set. By definition $A$ is NBB with respect to some linear order, notice that we can always choose this linear order to start with the atoms in $A$. We denote a linear order starting with $A$ and for which $A$ is $NBB$ by $\omega_{A}$. Let $A=B\sqcup C$ be a partition of $A$ that is NBB with respect to a linear order $\omega_{A}$ starting with the atoms in $B$. We denote such an order by $\omega_{B,C}$.
@@ 1. Restriction:
    Let $X\in\mathcal{L}$ be an element of rank $\geq 1$. The restriction of $\mathcal{L}$ to $X$ is the finite, atomic, and graded lattice $[\hat{0},X]$.
@@ 2. Contraction:
    Let $X\in\mathcal{L}$ be an element of rank $\leq rank(\mathcal{L})-1$. The contraction of $\mathcal{L}$ to $X$ is the finite and graded lattice $[X,\hat{1}]$.
@@ 3. Truncation:
    Let $m\leq rank(\mathcal{L})$. The truncation of $\mathcal{L}$ by $m$ is the finite, atomic, and graded lattice $T(\mathcal{L},m)=\left\{X\in\mathcal{L}\mid rank(X)<m\right\}\cup\left\{A(\mathcal{L})\right\}$.
@@ 4.2. Theorem:
The basic lattice operations above effect the family of independent sets of $\mathcal{L}$ in the following way:
@@
    $I([\hat{0},X])=\left\{I\cap A(X)\mid I\in I\left(\mathcal{L}\right)\right\}$ with $A(X)$ denoting the set of atoms $\leq X$.
@@
    $I([X,\hat{1}])\cong(\left\{I\mid\exists\omega_{I,B_{X}}\right\}\cap\left(\cup_{X\lessdot Y}A\left(Y\right)\right))/_{\sim_{X}}$ with $B_{X}$ a maximal independent set spanning $X$ and $\sim_{X}$ being the equivalence relation having $I\sim_{X}J$ if there exists $X\lessdot Y$ such that $I\setminus\left\{a\right\}=J\setminus\left\{b\right\}$ for some $a,b\in A(Y)\setminus A(X)$.
@@
    $I(T(\mathcal{L},m))=\left\{I\in I(\mathcal{L})\mid rank(I)\leq m\right\}$.
@@
    If $\mathcal{L}$ is coatomic, then a set of coatoms $\mathcal{X}\subset A(\mathcal{L}^{opp})$ is a geometric independent set iff there exist a linear order on $\mathcal{X}=\left\{X_{1},...,X_{n}\right\}$ and a geometric independent set $\left\{a_{1},...,a_{d-1}\right\}$ of $\mathcal{L}$ with $\omega_{\left\{a_{1},...,a_{n}\right\}}$ such that $\land_{i=1}^{t}X_{i}=\vee_{i=1}^{d-t}a_{i}$ for all $t\in[n]$.
@@
$(1)$ follows directly from the fact that any linear order of the restriction can be extended to a linear order of $\mathcal{L}$ and vice versa.
$(3)$ is also obvious as a rank $k$ set of atoms being independent depends only on the rank $\leq k$ elements of $\mathcal{L}$.
For $(2)$ we first observe that if $\omega_{I,B_{X}}$ exists, then $I$ cannot contain two different atoms $a,b\in A(Y)\setminus A(X)$ for $X\lessdot Y$. Therefore, every atom in $I$ corresponds to a unique atom of $[X,\hat{1}]$. Consequently, by abuse of notation, we have $\omega_{I}$ a linear order in which $I$ is $NBB$ in $[X,\hat{1}]$. In the other direction, we again have every atom $Y$ of $[X,\hat{1}]$ corresponding to a unique atom of $\mathcal{L}/_{\sim_{X}}$ by choosing any $a\in A(Y)\setminus A(X)$. Taking an order for which $I$ is $NBB$ in $[X,\hat{1}]$ we must have (again abusing notations) $\omega_{I,B_{X}}$ in $\mathcal{L}$ for any maximal independent set $B_{X}$ spanning $X$.
$(4)$ holds as a result of the discussion under definition [3.3]. Take such an order of $\mathcal{X}$ and choose any geometric independent set $\left\{a_{1},...,a_{n}\right\}$ with $\omega_{\left\{a_{1},...,a_{n}\right\}}$ that spans $\land_{X\in\mathcal{X}}X$. As $\mathcal{X}$ is geometric we have $rank(\land_{i=1}^{t+1}X_{i})+1=rank(\land_{i=1}^{t}X_{i})$ for every $t\in[k-1]$ and so we can find atoms $a_{n+1},...a_{d-1}$ as required by enlarging $\left\{a_{1},...,a_{n+t}\right\}$ to a spanning set of $\land_{i=1}^{t-1}X_{i}$ for each $t$. The other direction follows from:
$n=rank(\vee_{i=1}^{n}a_{i})=rank(\land_{i=1}^{n}X_{i})=\left|\mathcal{X}\right|$
Comparing Theorem [4.2] with the effect of these operations for geometric lattices (or simple matroids), we see that deletion and truncation work exactly the same. In contrast, contracting an element depends on the ordering of the independent sets, even if it is atomic. One consequence of this difference is that the family of finite, atomic and graded lattices with $I(\mathcal{L})$ a matroid is not closed under the operation of taking minors, which are lattices obtained from $\mathcal{L}$ by a sequence of contractions and deletions. For example, the lattice $\mathcal{L}=U_{6,4}\setminus\left\{\left\{123\right\},\left\{124\right\},\left\{126\right\},\left\{134\right\},\left\{136\right\}\right\}$, with $U_{6,4}$ being the geometric lattice of the uniform matroid of rank $4$ on the ground set $[6]$. One can easily check that $\mathcal{L}$ is an atomic graded lattice with $I(\mathcal{L})=I(U_{6,4})$ (obviously a matroid). However, the contraction $[1,[6]]$, which is still a finite, atomic and graded lattice, has a nonmatroid family of independent sets $I([1,[6]])=I(U_{5,3})\setminus\left\{\left\{234\right\},\left\{236\right\}\right\}$.
@@ 4.3. Definition:
Let $\mathcal{L}$ be a finite, atomic and graded lattice, a maximal independent set $I\in I(\mathcal{I)}$ will be called a basis of $\mathcal{L}$ if every subset of $I$ is geometric.
Our reason for calling these maximal independent sets bases is that, in the same way as bases of matroids, they admit the basis-exchange property. That is, if $B$ is a basis and $I$ is a maximal independent set, then for every atom $a\leq B$ there exists an atom $b\in I$ such that $(B\setminus\left\{a\right\})\cup\left\{b\right\}$ is a maximal independent set. This property comes from the fact that $\vee(B\setminus\left\{a\right\})\neq\hat{1}$ for every $a\in B$.
Another matroid-like property bases possess is the existence of a fundamental circuit.
@@ 4.4. Lemma:
Let $B$ be a basis for $\mathcal{L}$. For every $a\notin B$, $B\cup\left\{a\right\}$ contains a unique minimal dependent set denoted $\mathcal{C}_{B,a}$
@@
Assume $C_{1},C_{2}$ are minimal dependent sets contained in $B\cup\left\{a\right\}$ and let $x\in C_{1}\setminus C_{2}$. As $C_{1}\setminus\left\{x\right\}$ is independent we have $\vee(C_{1}\setminus\left\{x\right\})=\vee(C_{1}\setminus\left\{a\right\})$ and therefore a geometric independent set. We then have $(B\cup\left\{a\right\})\setminus\left\{x\right\}$ independent since adding each atom in $B\setminus\left\{x\right\}$ to $C_{1}\setminus\left\{x\right\}$ (in any order) will increase its rank. We now have $C_{2}\subseteq(B\cup\left\{a\right\})\setminus\left\{x\right\}$, which contradicts that it is dependent.
We continue our discussion of bases by describing the image of bases under basic lattice operations. From Theorem [4.2] we can see that truncating or taking minors preserve the bases of the lattice. To see that the same is true for dual operation, we observe that each subset $I^{\prime}$ of size $rank(\mathcal{L})-1$ of a basis $\mathcal{I}$ corresponds to the unique coatom $\vee I^{\prime}$. We define $I^{opp}=\left\{\vee(I\setminus\left\{a\right\})\mid a\in I\right\}$ as the basis of $\mathcal{L}^{opp}$ corresponding to $I$. $\mathcal{I}^{opp}$ is a basis, as for every subset $\mathcal{J}\subsetneq\mathcal{I}^{opp}$ and every $I^{opp}\in\mathcal{I}^{opp}\setminus\mathcal{J}$ we have $rank(\vee_{\mathcal{L}^{opp}}\mathcal{J})\lneq rank(\vee_{\mathcal{L}^{opp}}(\mathcal{J}\cup\left\{I^{opp}\right\}))$.
@@ 4.5. Theorem:
Let $I,J$ be bases of $\mathcal{L}$ with $opp(I)=opp(J)$. For all $a\in I$ there exists a unique $b\in J$ with $\vee(I\setminus\left\{a\right\})=\vee(J\setminus\left\{b\right\})$. $opp$ being injective now follows by induction on $I\setminus\left\{a\right\}$ and $J\setminus\left\{b\right\}$ as bases of $[\hat{0},\vee(I\setminus\left\{a\right\})]$.
Let $\mathcal{J}$ be a basis of $\mathcal{L}^{opp}$, for every $J\in\mathcal{J}$ we have a unique atom $a=\vee_{\mathcal{L}^{opp}}(\mathcal{J}\setminus\left\{J\right\})\in A(\mathcal{L})$ with $a\lneq J$ and $a\leq I$ for every $I\in\mathcal{J}\setminus\left\{J\right\}$. Therefore, $I=\left\{\vee_{\mathcal{L}^{opp}}(\mathcal{J}\setminus\left\{J\right\})\mid J\in\mathcal{J}\right\}$ is a basis of $\mathcal{L}$ such that $opp(I)=J$.
@@ 5. Adjoint matroids
In this section we apply the construction of the previous sections to the subject of adjoint matroids. We start by defining an important type of dependent sets related to $(\mathcal{L}^{*})^{opp}$, the dual lattice of the dual matroid. We have already shown that the size of an independent set of a lattice is $\leq$ its rank, therefore, every set larger than its rank must be dependent. We denote the complement of this family of sets by:
$\mathcal{S}(\mathcal{M})=\left\{\mathcal{S}\in 2^{\mathcal{C}(\mathcal{M})}\mid\left|\mathcal{S}\right|\leq rank_{(\mathcal{L}^{*})^{opp}}\left(\cap_{C\in\mathcal{S}}C^{c}\right)\right\}$
It was proven (for example, in "On Adjoints and Dual Matroids") that not all matroids have an adjoint and even if they have one it may not be minimal. We can reformulate Theorem [3.7] to give a necessary and sufficient condition for a matroid to be an adjoint of another matroid:
@@ 5.1. Theorem:
Let $\mathcal{M}$ be a matroid of rank $d$ on the ground set $[n]$ and $\mathcal{M}^{\triangle}$ a matroid of rank $n-d$ on the ground set $\mathcal{C}(\mathcal{M})$. $\mathcal{M}^{\triangle}$ is an adjoint of $\mathcal{M}$ iff $I((\mathcal{L}^{*})^{opp})\subseteq I(\mathcal{M}^{\triangle})\subseteq\mathcal{S}(\mathcal{M)}$.

### v3

@@ 4. Lattice operations
Let $\mathcal{L}$ be a finite, atomic, and graded lattice, and let $A\in I(\mathcal{L})$ be an independent set. By definition, $A$ is $NBB$ with respect to some linear order. Notably, we can always choose a linear order that begins with the atoms in $A$. We denote such an order, where $A$ remains $NBB$, by $\omega_{A}$. Now, let $B\subsetneq A$ such that $A$ is $NBB$ with respect to a linear order $\omega_{A}$ that starts with the atoms in $B$. We denote this order by $\omega_{A,B}$.
@@ 1. Restriction:
    Let $X\in\mathcal{L}$ be an element of rank $\geq 1$. The restriction of $\mathcal{L}$ to $X$ is the finite, atomic, and graded lattice $\hat{0},X]$. For clarity, we define $A(X)\coloneqq A([\hat{0},X])$.
@@ 2. Contraction:
    Let $X\in\mathcal{L}$ be an element of rank $\leq\rho(\mathcal{L})-1$. The contraction of $\mathcal{L}$ to $X$ is the finite and graded lattice $X,\hat{1}$].
@@ 3. Truncation:
    Let $m\leq\rho(\mathcal{L})$. The truncation of $\mathcal{L}$ by $m$ is the finite, atomic, and graded lattice $T(\mathcal{L},m)=\left\{X\in\mathcal{L}\mid\rho(X)<m\right\}\cup\left\{\hat{1}_{\mathcal{L}}\right\}$. To see the resulting poset is indeed a lattice observe that a lower set of a finite lattice is a finite meet semilattice, so that adding a top element makes it a lattice.
@@ 4. Dual:
4.2. Definition:
Let $\mathcal{L}$ be a finite, atomic and graded lattice. We say $\mathcal{L}^{\prime}$ is a minor of $\mathcal{L}$ if it is obtained from $\mathcal{L}$ by a sequence of contractions and restriction.
@@ 4.3. Theorem:
The basic lattice operations described above influence the family of independent sets of $\mathcal{L}$ in the following manner:
@@
    $I([\hat{0},X])=\left\{I\cap A(X)\mid I\in I\left(\mathcal{L}\right)\right\}$.
@@
    $I([X,\hat{1}])\cong(\left\{I\mid\exists\omega_{I\cup I_{X},I},\vee I_{X}=X\right\}\cap\left(\cup_{X\lessdot Y}A\left(Y\right)\right))/_{\sim_{X}}$, with $\sim_{X}$ being the equivalence relation on $A(\mathcal{L})$ having $a\sim_{X}b$ if there exists $X\lessdot Y$ such that $a,b\in Y\setminus X$. The first set can be best understood by selecting an independent set $I_{x}$ that spans $X$, then examining all $NBB$ that include $I_{X}$ while following a linear order that begins with the atoms outside $I_{X}$. Next, we remove atoms that do not belong to $X,\hat{1}$] and quotient by atoms that are smaller than the same covering element.
@@
    $I(T(\mathcal{L},m))=$
    $\left\{I\in I(\mathcal{L})\mid\rho(I)\leq m\right\}\cup\left\{I\in I(\mathcal{L})\mid\exists a\in I\text{ such that }\rho(I\setminus\left\{a\right\})\lneq m\right\}$.
@@
    If $\mathcal{L}$ is coatomic, then a set of coatoms $\mathcal{X}\subset A(\mathcal{L}^{opp})$ is a geometric independent set if and only if there exist a linear order on $\mathcal{X}=\left\{X_{1},...,X_{n}\right\}$ and a geometric independent set $\left\{a_{1},...,a_{d-1}\right\}$ of $\mathcal{L}$ with $\omega_{\left\{a_{1},...,a_{n}\right\}}$ such that $\land_{i=1}^{t}X_{i}=\vee_{i=1}^{d-t}a_{i}$ for all $t\in[n]$.
@@
$(1)$ Given a linear order $\omega$ on $A(\mathcal{L})$, we can derive a linear order on the restriction by considering the induced order $\omega\mid_{X}$ on $A(X)$. An $NBB$ set $I$ in $\omega$ then induces an $NBB$ set $I\cap A(X)$ in $\omega\mid_{X}$. Conversely, a linear order on $A(X)$ can be extended to a linear order on $\mathcal{L}$ by appending the remaining atoms in any order at the end. In this induced order, every $NBB$ set of $\hat{0},X]$ remains an $NBB$ set of $\mathcal{L}$
[…]
$(3)$ Given a linear order $\omega$ on $A(\mathcal{L})=A(T(\mathcal{L},m))$, every $NBB$ set in $\mathcal{L}$ with rank $\leq m$ remains an $NBB$ set in $T(\mathcal{L},m)$. If $\rho(I)\gneq m$ and there exists $a\in I$ such that $\rho(I\setminus\left\{a\right\})\lneq m$, then there exists a linear order $\omega_{\left\{a\right\}}$ for which $I$ is $NBB$ in $\mathcal{L}$, and $I$ remains an $NBB$ set in $T(\mathcal{L},m)$ with respect to $\omega_{\left\{a\right\}}$. For the remaining case, where $\rho\gneq m$ and no $a\in I$ satisfies $\rho(I\setminus\left\{a\right\})\lneq m$, any linear order on $A(\mathcal{L})$ can be taken. Letting $a$ be the first atom, we have $a\leq_{\mathcal{L}}\vee(I\setminus\left\{a\right\})$. Thus, $I$ is not $NBB$ with respect to any order in the truncation.
$(4)$ holds as a result of Lemma 3.4. Consider such an order on $\mathcal{X}$ and select any geometric independent set $\left\{a_{1},...,a_{n}\right\}$ with $\omega_{\left\{a_{1},...,a_{n}\right\}}$ that spans $\land_{X\in\mathcal{X}}X$. Since $\mathcal{X}$ is geometric, we have $\rho(\land_{i=1}^{t+1}X_{i})+1=\rho(\land_{i=1}^{t}X_{i})$ for every $t\in[k-1]$. Thus, we can find atoms $a_{n+1},...a_{d-1}$ extending $\left\{a_{1},...,a_{n}\right\}$ to a spanning set of $\land_{i=1}^{t-1}X_{i}$ for each $t$. The other direction follows from:
$n=\rho(\vee_{i=1}^{n}a_{i})=\rho(\land_{i=1}^{n}X_{i})=\left|\mathcal{X}\right|$
Remark:
We compare the effect of contraction for geometric lattices (or simple matroids) with general finite, graded and atomic lattices. Contracting an element of a non-geometric lattice depends on the ordering of the independent sets, even if the contraction is atomic. One consequence of this difference is that the family of finite, atomic and graded lattices with $I(\mathcal{L})$ a matroid is not closed under the operation of taking minors. For example, the lattice $\mathcal{L}=\mathcal{L}(\mathcal{U}_{6,4})\setminus\left\{\left\{123\right\},\left\{124\right\},\left\{126\right\},\left\{134\right\},\left\{136\right\}\right\}$, with $\mathcal{U}_{6,4}$ being the uniform matroid of rank $4$ on the ground set $[6]$. One can easily check that $\mathcal{L}$ is an atomic and graded lattice with $I(\mathcal{L})=I(\mathcal{U}_{6,4})$ (obviously a matroid). However, the contraction $1,[6]]_{\mathcal{L}}$, which is still a finite, atomic and graded lattice, has a non-matroid family of independent sets $I([1,[6]]_{\mathcal{L}})=I(\mathcal{U}_{5,3})\setminus\left\{\left\{234\right\},\left\{236\right\}\right\}$.
@@ 4.4. Definition:
Let $\mathcal{L}$ be a finite, atomic and graded lattice, a maximal independent set $I\in I(\mathcal{L)}$ will be called a basis of $\mathcal{L}$ if every subset of $I$ is geometric.
Our reason for calling these maximal independent sets bases is that, in the same way as bases of matroids, they admit the following basis-exchange property.
4.5. Theorem:
Let $\mathcal{L}$ be a finite, atomic and graded lattice. If $B$ is a basis of $\mathcal{L}$ and $I$ is a maximal independent set, then for every atom $a\in B$ there exists an atom $b\in I$ such that $(B\setminus\left\{a\right\})\cup\left\{b\right\}$ is a maximal independent set.
This property comes from the fact that $\vee(B\setminus\left\{a\right\})\neq\hat{1}$ for every $a\in B$. Another matroid-like property bases possess is the existence of a fundamental circuit.
@@ 4.6. Lemma:
Let $B$ be a basis for $\mathcal{L}$. For every $a\notin B$, $B\cup\left\{a\right\}$ contains a unique minimal dependent set denoted $\mathcal{C}_{B,a}$ .
@@
Assume $C_{1},C_{2}$ are minimal dependent sets contained in $B\cup\left\{a\right\}$ and let $x\in C_{1}\setminus C_{2}$. $C_{1}\setminus\left\{x\right\}$ and $C_{1}\setminus\left\{a\right\}$ are independent, adding an atom not in the span of an independent set to it is again independent. Therefore, $\vee C_{1}=\vee(C_{1}\setminus\left\{x\right\})=\vee(C_{1}\setminus\left\{a\right\})$ and therefore a geometric independent set. We then have $(B\cup\left\{a\right\})\setminus\left\{x\right\}$ independent since adding each atom in $B\setminus\left\{x\right\}$ to $C_{1}\setminus\left\{x\right\}$ (in any order) will increase its rank, and as $\vee(C_{1}\setminus\left\{x\right\})=\vee(C_{1}\setminus\left\{a\right\})$ this is also true for $C_{1}\setminus\left\{a\right\}$. We now have $C_{2}\subseteq(B\cup\left\{a\right\})\setminus\left\{x\right\}$, which contradicts $C_{2}$ being dependent.
We continue our exploration of bases by examining how they transform under the operation of taking duals. Our goal is to define the dual "image" of bases, we observe that each subset $I^{\prime}$ of size $\rho(\mathcal{L})-1$ of a basis $\mathcal{I}$ corresponds to the unique coatom $\vee I^{\prime}$. We define $I^{opp}=\left\{\vee(I\setminus\left\{a\right\})\mid a\in I\right\}$ as the basis of $\mathcal{L}^{opp}$ corresponding to $I$. To see $\mathcal{I}^{opp}$ is a basis, observe that for every subset $\mathcal{J}\subsetneq\mathcal{I}^{opp}$ and every $I^{opp}\in\mathcal{I}^{opp}\setminus\mathcal{J}$ we have $\rho(\vee_{\mathcal{L}^{opp}}\mathcal{J})\lneq\rho(\vee_{\mathcal{L}^{opp}}(\mathcal{J}\cup\left\{I^{opp}\right\}))$.
@@ 4.7. Theorem:
Let $I,J$ be bases of $\mathcal{L}$ with $opp(I)=opp(J)$. For all $a\in I$ there exists some $b\in J$ with $\vee(I\setminus\left\{a\right\})=\vee(J\setminus\left\{b\right\})$. oppopp being injective now follows by induction on $I\setminus\left\{a\right\}$ and $J\setminus\left\{b\right\}$ as bases of $\hat{0},\vee(I\setminus\left\{a\right\})]$.
Let $\mathcal{J}$ be a basis of $\mathcal{L}^{opp}$, for every $J\in\mathcal{J}$ we have a unique atom $a=\vee_{\mathcal{L}^{opp}}(\mathcal{J}\setminus\left\{J\right\})\in A(\mathcal{L})$. Therefore, $a\nleq J$ and $a\leq I$ for every $I\in\mathcal{J}\setminus\left\{J\right\}$. Consequently, $I=\left\{\vee_{\mathcal{L}^{opp}}(\mathcal{J}\setminus\left\{J\right\})\mid J\in\mathcal{J}\right\}$ is a basis of $\mathcal{L}$ such that $opp(I)=J$.
@@ 5. Adjoint matroids
In this section, we extend the construction developed in previous sections to the study of adjoint matroids. We begin by recalling a well-established result that links the coatoms—also referred to as hyperplanes—of a geometric lattice with the circuits of its dual, as presented in [1].
5.1. Lemma:
Let $\mathcal{M}$ be a matroid and $\mathcal{M}^{*}$ its dual matroid. The circuits of $\mathcal{M}^{*}$ are exactly the complement of the coatoms of $\mathcal{L}(\mathcal{M})$.
We continue with a definition of an important type of dependent sets related to $(\mathcal{L}^{*})^{opp}\coloneqq(\mathcal{L}(\mathcal{M})^{*})^{opp}$, the dual lattice of the dual matroid. We have already shown in Lemma 3.4 that the size of an independent set of a lattice is $\leq$ its rank, therefore, every set larger than its rank must be dependent. We denote the complement of this family of sets by:
$\mathcal{S}(\mathcal{M})=\left\{\mathcal{S}\in 2^{\mathcal{C}(\mathcal{M})}\mid\left|\mathcal{S}\right|\leq\rho_{(\mathcal{L}^{*})^{opp}}\left(\cap_{C\in\mathcal{S}}C^{c}\right)\right\}$
This definition is purposely written in a complicated way, to show that it depends only on the matroid $\mathcal{M}$. Thinking of circuits of $(\mathcal{M})$ as atoms of $(\mathcal{L}^{*})^{opp}$, and forgetting its complicated connection to $\mathcal{M}$, we are describing sets of atoms not larger than their rank.
Example:
Let $\mathcal{U}_{3,3}$ be the uniform matroid of rank $3$ on the ground set $[6]$. Its circuits are all the subsets of $[6]$ of size $4$ and therefore the atoms of $(\mathcal{L}(\mathcal{U}_{3,3})^{*})^{opp}$ are the sets of size $2$. We already know that pairs of atoms are always independent, so every pair is in $\mathcal{S}(\mathcal{U}_{3,3})$. We have $\rho((\mathcal{L}(\mathcal{U}_{3,3})^{*})^{opp})=3$, and so sets of four atoms are never in $\mathcal{S}({\mathcal{U}_{3,3}})$. Sets of three atoms $\left\{a,b,c\right\}$ are in $\mathcal{S}({\mathcal{U}_{3,3}})$ if and only if $\left|a\cup b\cup c\right|\geq 4$.
It was proven (for example, in "On Adjoints and Dual Matroids") that not all matroids have an adjoint and even if they have one it may not be minimal. We can reformulate Theorem 3.8 to give a necessary and sufficient condition for a matroid to be an adjoint of another matroid:
@@ 5.2. Theorem:
Let $\mathcal{M}$ be a matroid of rank $d$ on the ground set $[n]$ and $\mathcal{M}^{\triangle}$ a matroid of rank ndn-d on the ground set $\mathcal{C}(\mathcal{M})$. $\mathcal{M}^{\triangle}$ is an adjoint of $\mathcal{M}$ if and only if $I((\mathcal{L}^{*})^{opp})\subseteq I(\mathcal{M}^{\triangle})\subseteq\mathcal{S}(\mathcal{M)}$.